# Dynamic Reconstruction of Hand-Object Interaction with Distributed Force-aware Contact Representation

### v1

@@ 4.3 Force-aware Hand-pose Optimization, FO
Given the contact energy defined in Sec. 3, we optimize the predicted pose $\theta$ from the hand decoder, with respect to the reconstructed object mesh, to obtain a better hand mesh and contact state. Based on Equ. 3, we use the ball-query method to find the corresponding object vertices for each hand region with a radius $R$, and the point pair will be set up between them and the hand keypoints $\mathcal{J}_{h}$.
@@
${}^{*}\theta=\mathop{\arg\min}_{\theta}(E+\mathcal{L}_{r}+\mathcal{L}_{o})$ (10)
@@
5 Hand-Object Tactile Dataset, HOT Dataset
Since obtaining the ground truth in the real world, such as object deformation, is relatively hard, previous works [3, 33] on hand-object interaction recording mainly focus on rigid or articulated objects. To fully benchmark our method’s ability on deformable objects, we also create Hand-Object-Tactile dataset, HOT dataset, which contains depth images, force sensor readings, and SDF ground truths generated by the simulation environment ZeMa [8] which realizes intersection-free and high accuracy of contact modeling and deformation based on FEM.
5.1 Data Acquisition
An example of collecting a training data sequence in the simulation environment is shown in Fig. 3.b. To ensure valid data points and stable grasps with multiple contacts between the objects and various regions of the hand, we use DiPGrasp [30], a fast and effective grasp planner. After obtaining a feasible grasp pose, we fix the selected object model, position the hand a certain distance from the final grasp translation, and add a slight perturbation to the final rotation. The hand is set fully open with joint states their lower limits. In the first stage, the hand moves slowly towards the final grasp transformation target. Next, we perform the grasp by gradually increasing the joint state values until they reach the grasp joint target. This process yields a complete training data sequence, including depth images, point clouds, and tactile arrays $\{\mathcal{P}_{t},\mathcal{M}_{t}\}_{t=1}^{n}$ for each frame.
Figure 3: (a): The 22 regions and a typical distributed tactile sensor layout [24]. (b): An example of collecting data in the simulation environment. For every frame, it generates depth images collected by different views of cameras and tactile sensor readings computed by ZeMa.
5.2 Dataset Statistics
In the HOT Dataset, we adopt objects from YCB [1] model repositories, which contain various object categories. We selected 5 to 10 different objects in the category Bottle and Box from the YCB dataset, and we manually added 5 Sponges, 5 Plasticines, and 5 Stuffed Animals to demonstrate the ability of our model to track objects with large deformations.
The simulator ZeMa mainly uses density $\rho$, Young’s modulus $E$, and Poisson’s ratio $\nu$ to describe the properties of different deformable objects. We will therefore use the parameters of objects that correspond to reality. For the object in category Bottle and Box, we set $\rho=10^{3}kg/m^{3}$, $E=1.271$GPa, $\nu=0.28$, and for the Sponges, Plasticines and Stuffed Animals, $\rho=30kg/m^{3}$, $E=0.1$MPa, $\nu=0.38$. And as we consider Plasticine as plastic objects, we set its yield stress $s_{y}=200$pa.
For each object, we randomly generate 20 different trajectories, with 15 for training, 2 for testing, and 3 for validation and visualization. The whole dataset consists of 600 sequences which contains 50-100 frames, and each sequence involves 8 different views of camera for capturing point clouds.
6 Experiments
6.1 Implementation Details
We set the input point cloud size to $N=1024$. We randomly sample $1\times 10^{6}$ positions in the space, including $2\times 10^{5}$ points on the object surface. By subsampling to $M=2048$, we obtain the sample positions for SDF prediction during training. The point cloud feature size is set to $d=128$, the disparity feature size to $d^{\prime}=64$, and the resolution of the volume feature for SDF decoding to $D=64$. We empirically set the loss weights to $\lambda_{f}=0.01$, $\lambda_{S}=0.5$, and $\lambda_{H}=1$.
We first train our VDT-Net on the entire dataset with a batch size of 6 and a learning rate of $41e-$ 4 using the Adam optimizer for 100 epochs. We then fine-tune each object category with a batch size of 4 and a learning rate of $55e-$ 5 for 50-100 epochs, depending on the number of objects. The training process takes 15 hours on an Nvidia A40 GPU.
For the FO process, we set the threshold distance in the barrier function to $\hat{l}=2$mm and the ball-query radius for finding object vertices to $R=5$mm. We minimize the energy function for 100 iterations per frame with a learning rate of $32e-$ 3 using the Adam solver. It takes approximately $3.5\pm 0.5$ seconds per frame to complete refinement.
6.2 Datasets
To train and evaluate our method, we first use the dataset DexYCB, which captures hand grasping of rigid objects, with RGB-D cameras recorded. It can directly validate our method’s capability against the baselines. In addition, we use the HOT dataset, which contains interactions of hand and deformable objects with rich and accurate tactile information. It can evaluate the ability of our method to track object deformation and the effectiveness of our designed force-aware contact representation.
6.3 Metrics
Intersection over Union (IoU) evaluates the intersection percentage between the predicted object mesh and the ground truth.
Chamfer Distance (CD) assesses the reconstruction accuracy of object vertices.
Mean Per Joint Position Error (MPJPE) evaluates the accuracy of hand joint location tracking.
Penetration Depth (PD) measures the physical plausibility of hand-object interaction by reporting the maximum distance of hand penetration into the object.
Hand-Object Contact Mask IoU (CIoU) validates the contact recovery of hand-pose optimization by measuring the IoU of the contact mask between the optimized hand-object state and the ground truth. Contacts are defined as distances less than 3mm between hand and object vertices.
6.4 Results
For hand-object tracking evaluation, we compare our results with the state-of-the-art methods gSDF [6] and HOTrack [4]. gSDF uses RGB images to reconstruct objects with SDF representation, while HOTrack uses segmented point clouds to predict object poses. Our method uses unsegmented point clouds as input and output object meshes, making it more challenging to achieve better results. Additionally, since these baselines only work with rigid objects, we compare them only with the DexYCB object categories. DexYCB does not provide tactile information, so we only use VDT-Net to compare with the baseline methods here. Later, we provide an empirical way to use FO with a fixed force setting, which will be discussed in Sec. 6.5.
The quantitative results are shown in Tab. 1. In the DexYCB dataset, our leading scores in IoU and Chamfer distance demonstrate superior object tracking and reconstruction capabilities. Additionally, our higher MPJPE score indicates better hand-tracking performance. gSDF’s better performance on penetration depth is due to occasional incomplete object reconstruction, as illustrated in Fig. 4, where the pitcher is missing its handle. For the HOT dataset, VDT-Net achieves excellent scores. Significant improvements in MPJPE, PD, and CIoU after applying our force-based optimization confirm the effectiveness of our design.
@@ 6.4 Results
gSDF 86.8 13.4 14.4 8.9 31.3
HOTrack 88.2 10.2 25.7 12.3 28.5
Ours (VDT) 90.1 9.6 13.2 9.9 35.4
@@
Ours (VDT) 80.5 10.9 13.6 10.7 29.8
Ours (VDT + FO) * * 11.3 7.3 40.3
@@
Table 1: Quantitative results on DexYCB and HOT datasets for previous SOTA and our method. We only compare our results on rigid objects with the baselines. * indicates that adding FO doesn’t impact the metrics on reconstructed objects. $\uparrow$/$\downarrow$ / ↓ indicates higher scores/lower scores are better.
Fig. 4 shows the qualitative results of our method. Our tracking and reconstruction outperform state-of-the-art methods in the DexYCB dataset and show excellent performance on deformable objects in the HOT dataset. Our method accurately tracks hand movements and object deformations, thanks to the fusion of flow features and original point cloud features using a transformer, which incorporates both the 3D information of the current frame and correspondence features from the previous frame.
Fig. 5 demonstrates the performances of force-aware hand-pose optimization. Due to the design of force-aware contact representation, the optimization can solve most of the penetration problems and refine the contact map based on collected tactile information.
@@
Figure 4: Qualitative results on both DexYCB and HOT datasets. The upper shows our better performances on rigid objects compared to the baselines, and the lower demonstrates the effectiveness of VDT-Net in dynamic reconstruction for deformable objects and hands.
@@
Figure 5: Qualitative results with FO. As iteration steps rise, the penetration problem decreases, and the contact map is closer to the ground truth, making the contact more reasonable.
@@ 6.5 Ablation Study
We will conduct an ablation study on the following aspects: First, we demonstrate the results of introducing the distributed tactile arrays in the VDT-Net. Second, we discuss different definitions of force representation based on heuristic annotations or tactile-aware forms. In the appendix, we will also discuss the different point pair establishment methods in our contact representation and the importance of the flow prediction module.
Tactile feature fusion in VDT-Net. To assess the impact of introducing distributed tactile arrays in VDT-Net, we first use a 3-layer MLP to encode the tactile features of each region. By estimating hand pose, we fuse these regional features to the sample points, adding the encoded tactile feature to the point-wise feature of each sample position. We train VDT-Net with fused tactile information on our HOT dataset, and the results are shown in Tab. 3. Quantitative results show no significant improvements, likely because the tactile data are much more sparse than the visual inputs, causing feature misalignment. Therefore, we implement DF-Field to convert force readings into contact states for hand-pose optimization.
Metrics IoU(%$\uparrow$ CD(mm)$\downarrow$
VDT 80.5 10.9
VDT with Tactile 81.2 11.5
Table 2: Quantitative results for whether or not fusing tactile information in VDT-Net.
Metrics MPJPE PD CIoU
@@
$\mathcal{M}$ fix) 12.9 8.5 36.8
VDT+FO 11.3 7.3 40.3
@@
Table 3: Quantitative results for different force representations.
Different force representations. We also discuss the effect of different force representations on our contact representation. For the DexYCB dataset, since no tactile information was provided, we test our hand-pose optimization process with the force reading $\mathcal{M}_{j}=0.5$ for every region $j$ and compare the result with VDT-Net only. For the HOT dataset, we also conduct FO with $\mathcal{M}_{j}=0.5$ for testing the influence of removing tactile information.
The results are shown in Tab. 3. We can see that on the DexYCB dataset, even with $\mathcal{M}$ fixed, the results are also improving in all aspects, which proves the great effectiveness of our force-aware contact representation. As for the results on the HOT dataset, we can conclude that introducing tactile information improves both penetration problems and contact recovery, mostly because the distributed tactile arrays provide accurate hand-object interaction forces, which is more reliable than setting force parameters empirically.
7 Conclusion
@@
In this work, we introduce ViTaM-D, a dynamic hand-object interaction reconstruction framework that integrates distributed tactile sensing with visual perception. Featuring the DF-Field representation and force-aware optimization (FO), our approach effectively captures fine contact details and object deformation, and outperforms visual-only methods. We also present the HOT Dataset for benchmarking deformable object manipulation. Evaluations show that ViTaM-D outperforms state-of-the-art methods on both rigid and deformable objects. Future work includes integrating ViTaM-D with real tactile sensors and applying it to robot imitation learning, dexterous manipulation and human-robot collaboration.
Appendix A Flow Prediction Result
To validate the accuracy of our proposed flow prediction module (FPM), we report the Chamfer distance error on the DexYCB and HOT datasets in Tab. 4. The relatively low Chamfer distance demonstrates the efficacy of our network. The slightly better results on the DexYCB dataset are likely due to the larger hand-object movements and more challenging object deformations in the HOT dataset.
We also present some qualitative results in Fig. 6. The top section compares our predicted flow added to the last frame’s point cloud with the ground truth of the current frame’s point cloud. The near overlap of the two point clouds indicates high prediction accuracy. The bottom section shows a sequence of our estimation results, illustrating our method’s ability to track hand movements and object deformations in whole sequences.
Dataset DexYCB HOT
Category Box Bottle Sponge Plasticine
CD(mm)$\downarrow$ 8.7 9.1 10.3 12.2
Table 4: Quantitative results for flow predictions in VDT-Net.
Figure 6: Qualitative results on flow prediction using FPM.
Appendix B Ablation Study, Extended
B.1 Importance of flow prediction module
To demonstrate the importance of our flow prediction module (FPM), we conducted an experiment by directly fusing the visual features extracted from the input point clouds using the transformer fusion layer. The quantitative results are shown in Tab. 5. Introducing FPM significantly improves all scores, validating our feature fusion approach that incorporates the 3D static information of the current frame and the correspondence feature from the flow.
[…]
Dataset DexYCB
VDT(w/o FPM) 84.3 15.7 17.1
VDT 90.1 9.6 13.2
Dataset HOT
VDT(w/o FPM) 64.7 20.2 16.3
VDT 80.5 10.9 13.6
Table 5: Quantitative results on DexYCB and HOT Dataset for whether using the flow prediction module (denoted "FPM" in the table). "w/o" indicates without.
B.2 Different point pair establishment strategies
This section discusses the influence of two point pair establishment strategies: using keypoints or all-hand vertices. When considering all hand vertices, we establish point pairs between them and nearby object vertices, treating the force exerted by the hand as the reading from the nearest sensor. The quantitative results of these two methods are shown in Tab. 6. While penetration depth improves slightly, both contact IoU and MPJPE decrease. This is likely because considering all sensors leads to conflicting optimization directions for the same joint, as sensors within the same regions may experience different contact situations. Additionally, the iteration time increases about tenfold compared to our setting.
[…]
Keypoint 11.3 7.3 40.3 3.5 $\pm$ 0.5
All Verts 14.5 6.9 25.6 37 $\pm$ 3
Table 6: Quantitative results on evaluating point pair establishment on key points and on all-hand vertices.

### v2

@@ 4.3 Force-aware Hand-pose Optimization
Given the contact energy defined in Sec. 3, we optimize the predicted pose $\theta$ from the hand decoder, with respect to the reconstructed object mesh, to obtain a better hand mesh and contact state. Based on Eq. 3, we use the ball-query method to find the corresponding object vertices for each hand region with a radius $R$, and the point pair will be set up between them and the hand keypoints $\mathcal{J}_{h}$.
@@
${}^{*}\theta=\mathop{\arg\min}_{\theta}(E+\mathcal{L}_{r}+\mathcal{L}_{o}).$ (10)
@@
gSDF [6] (RGB) 86.8 13.4 14.4 8.9 31.3
HOTrack [4] 88.2 10.2 25.7 12.3 28.5
Zhang et al. [40] 88.1 10.4 15.3 10.6 30.5
Ours (w/o Force Opt.) 90.1 9.6 13.2 9.9 35.4
@@
Zhang et al. [40] 77.9 16.8 15.6 11.9 26.8
ViTaM [15] 80.5 11.5 15.1 10.6 28.5
Ours (w/o Force Opt.) 81.0 10.9 13.6 10.7 29.8
Ours (w. CPF [33]) * * 11.9 7.8 37.5
Ours (w. TOCH [42]) * * 13.5 9.3 34.1
Ours (w. Force Opt.) * * 11.3 7.3 40.3
@@
Table 1: Quantitative results on DexYCB and HOT datasets for previous SOTA and our method. We only compare our results on rigid objects with the baselines. * indicates that adding hand-object refinement methods doesn’t impact the metrics on reconstructed objects. $\uparrow$/$\downarrow$ / ↓ indicates higher scores/lower scores are better. Force Opt. is short for force-aware optimization.
5 Hand-Object Tactile Dataset, HOT Dataset
Obtaining ground truth for object deformation in real-world scenarios is challenging, and prior works [3, 35] on hand-object interaction primarily focus on rigid or articulated objects. To benchmark our method’s performance on deformable objects, we introduce the Hand-Object-Tactile (HOT) dataset, featuring depth images, force sensor readings, and SDF ground truths generated using the ZeMa simulator [8]. ZeMa employs FEM for intersection-free, high-accuracy contact modeling and deformation.
5.1 Data Acquisition
An example of data collection in the simulation environment is shown in Fig. 2.b. To ensure stable grasp with multiple contacts, we use DiPGrasp [32], a fast grasp planner. After obtaining a feasible grasp pose, we position the hand near the target, add slight rotational perturbations, and set the hand fully open. It then moves toward the target and performs the grasp, generating a sequence of depth images, point clouds, and tactile arrays $\{\mathcal{P}_{t},\mathcal{M}_{t}\}_{t=1}^{n}$ per frame.
5.2 Dataset Statistics
The HOT dataset includes objects from the YCB repository [1], with 5-10 objects from the Bottle and Box categories. We also add 5 Sponges, 5 Plasticines, and 5 Stuffed Toys to evaluate large deformations. Object properties are defined using density $\rho$, Young’s modulus $E$, and Poisson’s ratio $\nu$. For Bottle and Box, $\rho=10^{3}kg/m^{3}$, $E=1.271$GPa, $\nu=0.28$; for Sponges, Plasticines, and Stuffed Toys, $\rho=30kg/m^{3}$, $E=0.1$MPa, $\nu=0.38$. Plasticine has a yield stress $s_{y}=200$pa.
Each object has 20 trajectories (15 training, 2 testing, 3 validation), totaling 600 sequences of 50-100 frames. Each sequence includes point clouds from 8 camera views.
6 Experiments
@@
Figure 4: Qualitative results on both DexYCB and HOT datasets. The upper shows our better performances on rigid objects compared to the baselines, and the lower demonstrates the effectiveness of model for dynamic tracking of deformable objects and hands.
6.1 Implementation Details
We use Farthest Point Sampling to downsample the input point cloud size to $N=1024$ and sample $1\times 10^{6}$ positions in space, including $2\times 10^{5}$ surface points. During training, we subsample $M=2048$ for SDF prediction. The volume scattering feature resolution $D=64$. The distance threshold in contact constraint $l_{c}=3$mm. Loss weights are $\lambda_{f}=0.01$, $\lambda_{S}=0.5$, $\lambda_{H}=1$, and $\lambda_{C}=0.05$.
The visual dynamic tracking net is trained on the entire dataset with a batch size of 6, learning rate $41e-$ 4, and Adam optimizer for 100 epochs. Fine-tuning per object category uses a batch size of 4, learning rate $55e-$ 5, and 50-100 epochs. Training takes 15 hours on an Nvidia A40 GPU.
For force-aware optimization, the barrier function threshold $\hat{l}=2$mm and the ball-query radius $R=5$mm. The energy is minimized for 100 iterations per frame with a learning rate of $32e-$ 3, taking $3.5\pm 0.5$ seconds per frame.
6.2 Datasets
We use DexYCB for rigid object testing and the HOT dataset for deformable object interactions. DexYCB validates our method against baselines, while HOT evaluates deformation tracking and force-aware contact modeling.
6.3 Metrics
Intersection over Union (IoU) measures mesh overlap between predictions and ground truth.
Chamfer Distance (CD) evaluates object vertex reconstruction accuracy.
Mean Per Joint Position Error (MPJPE) assesses hand joint tracking precision.
Penetration Depth (PD) quantifies hand-object interaction plausibility by measuring maximum hand penetration into the object.
Hand-Object Contact Mask IoU (CIoU) validates contact recovery by comparing contact masks (defined as distances $<3$mm) between optimized results and ground truth.
6.4 Results
For hand-object tracking evaluation, we compare our method with state-of-the-art approaches: gSDF [6], HOTrack [4], Zhang et al. [40], ViTaM [15], CPF [33], and TOCH [42]. We include gSDF which uses RGB images only for metrics references, and [40] uses depths for SDF-based object tracking. HOTrack predicts object poses from segmented point clouds, and ViTaM, similar to our method, takes unsegmented point clouds and tactile arrays as input but uses WNF representation, making it a strong baseline. We compare gSDF and HOTrack on DexYCB for rigid objects without tactile input, and ViTaM on the HOT dataset due to its tactile requirement. CPF and TOCH are used to benchmark our force-based optimization. We will discuss how to use force-aware optimization empirically with a fixed force setting in Sec. 6.5.
Quantitative results in Tab. 1 show our method outperforms baselines on DexYCB in IoU and Chamfer distance, demonstrating superior object tracking and reconstruction. Our higher MPJPE score indicates better hand-tracking performance. On the HOT dataset, our visual dynamic tracking net matches ViTaM’s performance, while force-based optimization significantly improves MPJPE, PD, and CIoU, confirming its effectiveness. This is largely due to our method’s ability to obtain accurate contact information through tactile arrays, unlike previous methods that rely on network-based contact modeling.
@@ 6.4 Results
Figure 5: Qualitative results with force-aware optimization. As iteration steps rise, the penetration problem decreases, and the contact map is closer to the ground truth, making the contact more reasonable.
Figure 4 shows the qualitative results of our method. Our tracking and reconstruction outperform state-of-the-art methods in the DexYCB dataset, and our force-aware optimization process shows excellent performance on deformable objects in the HOT dataset in comparison with previous hand-object optimization baselines. Our method accurately tracks hand movements and object deformations, thanks to the fusion of flow features and original point cloud features using a transformer, which incorporates both the 3D information of the current frame and correspondence features from the previous frame. Figure 5 visualizes the force-aware hand-pose optimization. With force-aware contact representation DF-Field, the optimization can solve most of the penetration problems and refine the contact map based on collected tactile information.
@@ 6.5 Ablation Study
We conduct the ablation study on (1) methods for obtaining contact states for the contact constraint (Sec. 4.2) and (2) definitions of force representation (heuristic vs. tactile-aware). Additional discussions on point pair establishment, flow prediction, and tactile array integration in the visual dynamic tracking net are in the supplementary material.
Contact State & Effect of Contact Constraint. We evaluate the impact of the contact constraint by testing the visual dynamic tracking net without it. Additionally, we compare ground truth contact states with predictions from a PointNet-based network when there is a lack of tactile hardware support. Results in Tab. 2 show that the contact constraint significantly improves object reconstruction and hand-object interaction prediction. The PointNet-based network effectively predicts interaction states from visual inputs, while tactile readings match ground truth accuracy, greatly enhancing object geometry reconstruction.
[…]
w/o CC 75.9 12.8 25.3 12.0
CC. (GT) 81.2 11.2 29.2 11.9
CC. (Tactile) 81.0 10.9 29.8 11.9
CC. (PointNet) 78.3 12.1 27.6 12.1
Table 2: Quantitative results for different contact information acquirement. CC. indicates Contact Constraint, along with approaches for acquiring contact information.
[…]
@@
w/o Force Opt. 13.6 10.7 29.8
$\mathcal{M}$ fixed Force Opt. 12.9 8.5 36.8
Force Opt. 11.3 7.3 40.3
@@
Table 3: Quantitative results for different force representations in force-aware hand-pose optimization.
Different Force Representations. We explore the effect of force representations on contact modeling. For DexYCB (no tactile data), we set $\mathcal{M}_{j}=0.5$ for regions near objects ($l=3mm$) and compare results with the visual dynamic tracking net. For HOT, we test the optimization with $\mathcal{M}_{j}=0.5$ to evaluate the impact of removing tactile information.
Results in Tab. 3 demonstrate that even with fixed $\mathcal{M}$, DexYCB performance improves, validating the effective design of DF-Field. On HOT, tactile information significantly enhances penetration resolution and contact recovery, as distributed tactile arrays provide accurate interaction forces, outperforming empirical force settings.
6.6 Real-world Experiments
In order to validate the effectiveness of our method, we conduct real-world experiments with our network trained on simulation data. We bought a ViTaM glove [15] from its authors to capture distributed tactile sensory data and a Photoneo MotionCam M+ high-precision depth camera to record the point cloud series. We can see an example of a reconstruction result on an unseen stuffed toy in Fig. 6. We will demonstrate our real-world setup and more series of tracking results in supplementary material.
Figure 6: Real-world reconstruction example of an unseen stuffed toy. The overall shape and deformation of the toy have been well reconstructed.
7 Conclusion and Future Works
@@
In this work, we introduce ViTaM-D, a dynamic hand-object interaction reconstruction framework that integrates distributed tactile sensing with visual perception. Featuring DF-Field and force-aware optimization, our approach effectively captures fine contact details and object deformation. We also present the HOT Dataset for benchmarking deformable object manipulation. Evaluations show that ViTaM-D outperforms state-of-the-art methods on both rigid and deformable objects. Future works include applying ViTaM-D to data collection for robotic imitation learning, dexterous manipulation, and human-robot interaction.
Acknowledgement
This work was supported by the Shanghai Commitee of Science and Technology, China (Grant No.24511103200) by the National Key Research and Development Project of China (No.2022ZD0160102), Shanghai Artificial Intelligence Laboratory, XPLORER PRIZE grants. This work has also been supported by the Shanghai Municipal Education Commission (No. 2024AIYB010) and the Fundamental Research Funds for the Central Universities (YG2025LC03).